a colony of ants carried away 20 of your muffins. that was 5/6 of all of then! how many are left?
1
Expert's answer
2020-03-27T14:59:05-0400

Number of muffins carried by ants=20

Total number of muffins=X

20 is the 5/6 of total muffins, X

X*(5/6)=20

X=(20*6)/5=24

Total number of muffins is 24